The square of a positive number minus the original number equals 72. What is the original number?
6
7
8
9
10

Answer by oberobic(2304)

x = a positive number
The square of the positive number = x^2.
x^2 minus the original number = x^2 - x.
x^2 -x = 72 (Given)
.
Subtract 72 from both sides
x^2 -x - 72 = 0
Factor
(x - 9)(x + 8) = 0
So,
x = 9
or
x = -8
But, we're told it is a positive number, which mean
x = 9
Always check your answer!
x^2 - x = 81 - 9
x^2 - x = 72
True
Be sure to state your answer.
The original number is 9.
Done.